# Relative monodromy of ramified sections on abelian schemes

### v3

@@ 3.2 Non triviality of the relative monodromy
In this section we prove 1.2, i.e. that $M_{\sigma}^{\textnormal{rel}}$ is non-trivial.
We are going to combine some topological techniques and cohomology theory in order to get some control on the relative monodromy of sections. Let’s keep the same notation as above and denote by $T:=p(S)\subseteq\mathbb{A}_{g}$ the image of the modular map, where $\mathfrak{A}_{g}\to\mathbb{A}_{g}$ denotes some universal family of $g$-dimensional abelian varieties with some fixed $-\ell$-structure and without locally constant parts. Thanks to 3.1, we can assume that the abelian scheme $\phi:\mathcal{A}\to S$ is endowed with a generically finite modular map $p:S\to T$. In order to simplify the notation after eventually restricting to Zariski open dense subsets we can assume that $p:S\to T$ is finite. Let $\sigma:S\to\mathcal{A}$ be a ramified section.
@@
Notice that the notion of ramified section is not affected by the restriction of the base. In fact the ramification property must be checked after extending the section to a maximal base. One uses the fact that for any open dense subset $U\subseteq S$ that induces an abelian scheme $\mathcal{A}|_{U}\to U$ we have $\Sigma(S)\cong\Sigma(U)$.
One of the main ideas of the proof is to construct a certain “unramified family” (i.e. unramified base change of the universal family) associated to our abelian scheme, which forces some constraints on the logarithm of ramified sections. To this end, let’s consider the classifying map $p:S\to T$ and look at the induced homomorphism on fundamental groups, i.e. $p_{*}:\pi_{1}(S(\mathbb{C}),s)\to\pi_{1}(T(\mathbb{C}),x)$ where $s\in S(\mathbb{C})$ and $x\in T(\mathbb{C})$ are two base points such that $x=p(s)$. The inclusion $p_{*}\pi_{1}(S(\mathbb{C}),s)\subseteq\pi_{1}(T(\mathbb{C}),x)$ gives rise to a finite unramified morphism $q^{\prime}:S^{\prime}(\mathbb{C})\to T(\mathbb{C})$ that by [24, Corollary 12.19] satisfies
$q^{\prime}_{*}\pi_{1}(S^{\prime}(\mathbb{C}),s^{\prime})=p_{*}\pi_{1}(S(\mathbb{C}),s),$ (18)
where $s^{\prime}$ is a fixed base point with $q^{\prime}(s^{\prime})=x$. By Equation 18, we can lift $p$ to a morphism $q:S(\mathbb{C})\rightarrow S^{\prime}(\mathbb{C})$ such that $q(s)=s^{\prime}$. Considering the corresponding abelian schemes we obtain the diagram
${\mathcal{A}(\mathbb{C})}$${\mathcal{A}^{\prime}(\mathbb{C})}$${\mathfrak{A}_{g|_{T}}(\mathbb{C})}$${S(\mathbb{C})}$${S^{\prime}(\mathbb{C})}$${T(\mathbb{C}).}$$\scriptstyle{q}$$\scriptstyle{p}$$\scriptstyle{q^{\prime}}$ (19)
Notice that the abelian scheme $\mathcal{A}^{\prime}\to S^{\prime}$ is an unramified family by construction, since $q^{\prime}:S^{\prime}(\mathbb{C})\to T(\mathbb{C})\subseteq\mathbb{A}_{g}(\mathbb{C})$ is an unramified morphism.
@@
We want to point out some algebraic and topological properties of the above construction. First we show that $q$ and $q^{\prime}$ are both finite morphisms of schemes: since $p$ is finite, then $q$ and $q^{\prime}$ are quasi-finite. Since $q^{\prime}$ is a finite-sheeted (unramified) topological cover, then it is proper in the sense that the inverse image of every compact set is compact. This implies that $q^{\prime}$ is a proper morphism in the sense of scheme theory, and thus $q^{\prime}$ is finite. In particular, $q^{\prime}$ is also a separated morphism. Since $p=q^{\prime}\circ q$ is finite and $q^{\prime}$ is separated, then $q$ is proper. Since $q$ is proper and quasi-finite, then $q$ is finite.
@@
${\pi_{1}(S(\mathbb{C}),s)}$${\pi_{1}(S^{\prime}(\mathbb{C}),s^{\prime})}$${p_{*}\pi_{1}(S(\mathbb{C}),s),}$$\scriptstyle{q_{*}}$$\scriptstyle{p_{*}}$$\scriptstyle{q^{\prime}_{*}}$ (20)
where $q^{\prime}_{*}:\pi_{1}(S^{\prime}(\mathbb{C}),s^{\prime})\to p_{*}\pi_{1}(S(\mathbb{C}),s)$ is an isomorphism.
@@
In the opposite direction, since $q$ is a finite morphism, we define the trace operator
@@
$\operatorname{Tr}(\tau)(y^{\prime}):=\sum_{y\in q^{-1}(y^{\prime})}m_{y}\tau(y)\in\mathcal{A}^{\prime}_{y^{\prime}},$
@@
Remark 3.5.
When $q$ is generically finite (but not necessarily finite) we can anyway define a trace operator for the algebraic sections $\operatorname{Tr}:\Sigma(S)\to\Sigma^{\prime}(S^{\prime})$ restricting the previous construction to the open sets where $q$ is finite.
@@
$\operatorname{Tr}(\exp(f))=\exp^{\prime}(\operatorname{Tr}(f))\qquad\textrm{and}\qquad\exp(q^{\prime*}(f^{\prime}))=q^{\prime*}(\exp^{\prime}(f^{\prime})).$ (22)
@@
where $q_{C},q^{\prime}_{C^{\prime}}$ and $p_{C}$ denote the restrictions of the corresponding maps. In this situation, if $C^{\prime}$ and $D$ are still of dimension $1$, we can clearly consider the analogous notions of trace operator and pullback operator as well as the restriction of a section $\sigma:S\to\mathcal{A}$ to $C$, which will be denoted by $\sigma_{C}$. We also have analogous notions of monodromy representations and monodromy groups, e.g. we will write $\rho_{C},\theta_{\sigma_{C}},\textrm{Mon}(\mathcal{A}|_{C})$ and $M^{\textrm{rel}}_{\sigma_{C}}$ for the obvious objects associated to the scheme $\mathcal{A}|_{C}\to C$ and to the section $\sigma_{C}:C\to\mathcal{A}|_{C}$; we do the same for $C^{\prime}$. Moreover, if we denote by $i:C\hookrightarrow S(\mathbb{C})$ and $i^{\prime}:C^{\prime}\hookrightarrow S^{\prime}(\mathbb{C})$ the natural embeddings and choose base points $s\in C$ and $s^{\prime}\in C^{\prime}$, from a topological point of view we obtain the following diagram involving the induced homomorphisms between fundamental groups (for the surjectivity of the upper horizontal arrow see 3.4):
${\pi_{1}(S(\mathbb{C}),s)}$${\pi_{1}(S^{\prime}(\mathbb{C}),s^{\prime})}$${\pi_{1}(C,s)}$${\pi_{1}(C^{\prime},s^{\prime}).}$$\scriptstyle{q_{*}}$$\scriptstyle{q_{C*}}$$\scriptstyle{i_{*}}$$\scriptstyle{i^{\prime}_{*}}$ (24)
@@ 1. $(i)$
    the morphism $q_{C}:C\to C^{\prime}$ appearing in Equation 23 is finite, the curves $C$ and $C^{\prime}$ are affine and the abelian schemes $\mathcal{A}_{|_{C}}\to C$, $\mathcal{A}^{\prime}_{|_{C^{\prime}}}\to C^{\prime}$ have no fixed part;
@@ 2. $(ii)$
    $\operatorname{Tr}(\sigma_{C})=\operatorname{Tr}(\sigma)|_{C^{\prime}}$;
@@ 6. $(vi)$
    $\theta_{\sigma_{C}}=\theta_{\sigma}\circ i_{*}$ and $\theta_{\sigma}(\pi_{1}(S(\mathbb{C}),s))=\theta_{\sigma_{C}}(\pi_{1}(C,s))$;
@@ 7. $(vii)$
    if $M_{\sigma}^{\textrm{rel}}=\{0\}$, then $M_{\sigma_{C}}^{\textrm{rel}}=\{0\}$.
@@
    In order to construct the curve $C$ let us consider the general hyperplane section $Y=H\cap S(\mathbb{C})$ and choose a base point $s\in Y$. By Lefschetz theorem on quasi-projective varieties [19, 2.2], the natural map $\pi_{1}(Y,s)\to\pi_{1}(S(\mathbb{C}),s)$ between fundamental groups is surjective. After iterating the process of taking such generic hyperplane sections we obtain an irreducible affine curve $C_{1}$ with the property that the natural map $\pi_{1}(C_{1},s)\to\pi_{1}(S(\mathbb{C}),s)$ is surjective for a base point $s\in C_{1}$. The restriction of $q$ to $C_{1}$ defines a morphism $q_{C_{1}}$ which is finite: in fact, since every closed embedding is a finite morphism and $q:S\to S^{\prime}$ is a finite morphism (see 3.4), then the restriction $q_{C_{1}}$ is finite. We define $C^{\prime}:=q_{C_{1}}(C_{1})$ and $C:=q^{-1}(C^{\prime})$. Notice that $C^{\prime}$ is an affine curve since $q_{C_{1}}$ is finite and surjective; moreover, $C$ could be no longer irreducible but it is still affine because $q$ is an affine morphism (since it is finite). Furthermore, since $q:S\to S^{\prime}$ is finite then it is closed. We have a closed embedding $i:C\hookrightarrow S(\mathbb{C})$, and we can conclude that the restriction $q_{C}:C\to C^{\prime}$ is a finite morphism. The abelian scheme $\mathcal{A}_{|_{C}}\to C$ has no fixed part by genericity of the hyperplane sections, and this implies that also $\mathcal{A}^{\prime}_{|_{C^{\prime}}}\to C^{\prime}$ has no fixed part.
@@
    This property is true by construction, since $C=q^{-1}(C^{\prime})$.
@@
    By $(i)$ we have a surjective homomorphism $\pi_{1}(C_{1},s)\to\pi_{1}(S(\mathbb{C}),s)$. Since $C_{1}$ is an irreducible component of $C$ and since removing a finite number of points preserves the fact of having a surjective map on fundamental groups, then the morphism $i_{*}:\pi_{1}(C,s)\to\pi_{1}(S(\mathbb{C}),s)$ is surjective. Then, the surjectivity of $i^{\prime}_{*}:\pi_{1}(C^{\prime},s^{\prime})\to\pi_{1}(S^{\prime}(\mathbb{C}),s^{\prime})$ follows by Equation 24.
@@ 6. $(vi)$
    Since we have the embedding $i:C\hookrightarrow S(\mathbb{C})$, we clearly have $\theta_{\sigma_{C}}(\pi_{1}(C),s)\subseteq\theta_{\sigma}(\pi_{1}(S(\mathbb{C}),s))$. In the other direction, by $(iv)$ it is enough to show that each variation of a logarithm of $\sigma$ induced by an element of $\pi_{1}(S(\mathbb{C}),s)$ can be obtained as a variation of a logarithm of $\sigma_{C}$ induced by the monodromy action of some element of $\pi_{1}(C,s)$; and this is true thanks to the surjectivity of $i_{*}$ which was proven in part $(iii)$. This also proves the equality $\theta_{\sigma_{C}}=\theta_{\sigma}\circ i_{*}$.
@@ 7. $(vii)$
Let us study the exact sequence in cohomology groups induced by Equation 25. By 2.5 we can conclude that no non-zero period can be globally defined on $C$. In other words, we get $H^{0}(C,\Lambda_{C})=0$. In addition, since $C$ is a Stein space we have
@@
By 2.9, we know that the obstruction to define a global logarithm of a section is given by some Galois cohomology class. We want to explain the interplay between the relevant Galois cohomology group and the exact sequence in Equation 26. Let’s fix the notation $G:=\pi_{1}(C,s)$ for a base point $s\in C$ and denote by $\rho_{C}:G\to\textrm{Mon}(\mathcal{A}|_{C})\subseteq\textrm{Sp}_{2g}(\mathbb{Z})$ the monodromy representation associated to the abelian scheme $\mathcal{A}|_{C}\to C$; we use the notation $\overline{h}:=\rho_{C}(h)$. With the aim of proving how the Galois cohomology interacts with the sheaf cohomology in our situation, we put the following remark in order to interpret the sections of the sheaf of periods as continuous maps on the universal cover of $C$ which are well-behaved with respect to the action of $G$.
Remark 3.7.
Notice that we can view $\Lambda_{C}$ geometrically also as a covering space of the base $C$ with fibers isomorphic to $\mathbb{Z}^{2g}$. Denote by $u_{C}:\widetilde{C}\to C$ the universal cover of the curve $C$ and note that the pull-back of $\Lambda_{C}$ to the universal covering space $\widetilde{C}$ is the constant sheaf $\mathbb{Z}_{\widetilde{C}}^{2g}$. We can obtain the map $\Lambda_{C}\rightarrow C$ via the diagram
[…]
Observe that the group $G$ acts on $\widetilde{C}$ in the usual way and on $\mathbb{Z}^{2g}$ via its monodromy representation $\rho_{C}$; the space $C$ is then the orbit space of the diagonal action of $G$ on $\widetilde{C}\times\mathbb{Z}^{2g}$.
For every open set $V\subset C$, we can consider the open set $u_{C}^{-1}(V)$ which is invariant by the action of $G$ on $\widetilde{C}$. Thus, the group of sections $\Gamma(V,\Lambda_{C})$ corresponds to the group of continuous (i.e. locally constant) maps $w:u_{C}^{-1}(V)\rightarrow\mathbb{Z}^{2g}$ satisfying $\overline{h}w=w\circ h$, for all $h\in G$, i.e. $\overline{h}\cdot(w\circ h^{-1})=w$.
We now give a description of the group $H^{1}(C,\Lambda_{C})$ in terms of the Galois cohomology with respect to the action of $G$ on $\mathbb{Z}^{2g}$ induced by the projection $\rho_{C}$. We have the following
@@
Let $\mathcal{V}=(V_{i})_{i\in I}$ be an open covering of $C$ such that each open set $V_{i}$ and each non-empty intersection $V_{i,j}=V_{i}\cap V_{j}$ is contractible, so that the first cohomology space is determined by the cohomology classes of the cocycles associated to this covering. Let us define a homomorphism $H^{1}(\mathcal{V},\Lambda_{C})=H^{1}(C,\Lambda_{C})\rightarrow H^{1}(G,\mathbb{Z}^{2g})$ and prove it is an isomorphism.
Let $(\eta_{i,j})_{i,j}$ be a cocycle in $H^{1}(\mathcal{V},\Lambda_{C})$, where $\eta_{i,j}\in\Gamma(V_{i,j},\Lambda_{C})$. We refer to the diagram of 3.7 and denote by $u_{C}:\widetilde{C}\rightarrow C$ the universal cover of $C$; moreover, let us define $U_{i}=u_{C}^{-1}(V_{i})$ and $U_{i,j}=u_{C}^{-1}(V_{i,j})=U_{i}\cap U_{j}$. By considering the pull-backs $u_{C}^{*}(\eta_{i,j})$, we obtain a cocycle $w_{i,j}$ with values in $\mathbb{Z}^{2g}$ for the covering $u_{C}^{*}(\mathcal{V})=(\pi^{-1}(V_{i}))_{i\in I}$. By the above remark, the elements $w_{i,j}$ can be viewed as continuous (i.e. locally constant) functions $U_{i,j}\rightarrow\mathbb{Z}^{2g}$, satisfying
$w_{i,j}=\overline{h}\cdot(w_{i,j}\circ h^{-1})\qquad\textrm{for each }h\in G.$ (27)
Since $\widetilde{C}$ is simply connected, we can write
$w_{i,j}=w_{i}-w_{j}$
for suitable continuous functions $w_{i}:U_{i}\rightarrow\mathbb{Z}^{2g}$, for $i\in I$. Let us define, for every $h\in G$ and $i\in I$, the continuous function $\gamma_{h,i}:U_{i}\rightarrow\mathbb{Z}^{2g}$ by
$\gamma_{h,i}=w_{i}-\overline{h}\cdot(w_{i}\circ h^{-1}).$
[…]
$H^{1}(\mathcal{V},\Lambda_{C})\rightarrow H^{1}(G,\mathbb{Z}^{2g}),\qquad[(\eta_{i,j})_{i,j}]\mapsto[h\mapsto\gamma_{h}].$
Let us verify that this homomorphism is injective. Take then a cocycle $(\eta_{i,j})_{i,j}$ giving rise, via the above procedure, to a coboundary $h\mapsto\gamma_{h}$. Then we can write $\gamma_{h}=\gamma-\overline{h}\cdot\gamma$ for a fixed vector $\gamma\in\mathbb{Z}^{2g}$; now, setting $w^{\prime}_{i}=w_{i}-\gamma$ we obtain that
$w_{i,j}=w^{\prime}_{i}-w^{\prime}_{j},$
where the $w^{\prime}_{i}$ satisfy the invariance condition
$\overline{h}\cdot(w^{\prime}_{i}\circ h^{-1})=w^{\prime}_{i},\qquad\textrm{foreach }h\in G$
similar to the relation holding for the $w_{i,j}$. By 3.7, this fact means that the $w^{\prime}_{i}$ are of the form $w^{\prime}_{i}=u_{C}^{*}(\eta_{i})$ for suitable sections $\eta_{i}\in\Gamma(V_{i},\Lambda_{C})$, so that $\eta_{i,j}=\eta_{i}-\eta_{j}$ is a coboundary.
Let us now verify that the homomorphism $H^{1}(\mathcal{V},\Lambda_{C})\rightarrow H^{1}(G,\mathbb{Z}^{2g})$ is also surjective. Let then $h\mapsto\gamma_{h}$ be a cocycle with values in $\mathbb{Z}^{2g}$, so that it satisfies
$\gamma_{h_{1}h_{2}}=\overline{h}_{1}\cdot\gamma_{h_{2}}+\gamma_{h_{1}},\qquad\textrm{for each }h_{1},h_{2}\in G.$
We want to define functions $w_{i}:U_{i}\rightarrow\mathbb{Z}^{2}$ satisfying
$w_{i}-\overline{h}\cdot(w_{i}\circ h^{-1})=\gamma_{h}\qquad\textrm{for each }h\in G.$ (28)
Recall that each open set $U_{i}=p^{-1}(V_{i})$ is a disjoint union of connected open sets of the form $h(U^{0}_{i})$, for $h\in G$ and for some component $U^{0}_{i}$ of $U_{i}$ (actually any choice of a component would work). We can then choose the function $w_{i}:U_{i}\rightarrow\mathbb{Z}^{2g}$ so that it vanishes on $U^{0}_{i}$ and, on the component $h^{-1}(U^{0}_{i})$ its value is $-\overline{h}^{-1}\cdot\gamma_{h}$. Then the cocycle condition satisfied by $\gamma$ implies the relations Equation 28 on each component $h(U_{i}^{0})$, so on the whole $U_{i}$. At this point we have that the functions $w_{i}-w_{j}=:w_{i,j}$ on $U_{i}\cap U_{j}$ are invariant, in the sense that $\overline{h}\cdot(w_{i,j}\circ h^{-1})=w_{i,j}$, so they are of the form $u_{C}^{*}(\eta_{i,j})$ for sections $\eta_{i,j}\in\Gamma(V_{i,j},\Lambda_{C})$.
Therefore, thanks to the exact sequence Equation 26 and 3.8 we obtain a surjective group homomorphism
@@
Notice that by 3.6 also the curve $C^{\prime}$ is a Stein space. Moreover, by 2.5 we still obtain $H^{0}(C^{\prime},\Lambda_{C^{\prime}})=0$. Therefore, the analogous results of Equation 26 and 3.8 hold if we replace the curve $C$ by the curve $C^{\prime}$. In particular, for some base point $s^{\prime}\in C^{\prime}$ we have the surjective group homomorphism

### v4

@@ 4.2 Non triviality of the relative monodromy
In this section we prove 1.2, i.e. that $M_{\sigma}^{\textnormal{rel}}$ is non-trivial for all sections which ramify outside a fixed proper closed subset of the base. Recall that by 4.1, we can assume that the abelian scheme $\phi:\mathcal{A}\to S$ is endowed with a dominant morphism $p^{\nu}:S\to T^{\nu}$ coming from the normalization of $T$.
By the hypothesis of the theorem, the modular map $p\colon S\to T$ is assumed to be generically finite. Using [28, Lemma 10.5], this implies that the locus $U_{p}\subseteq S$ where $p$ is finite and flat forms a dense Zariski-open subset. The non-triviality of $M_{\sigma}^{\textnormal{rel}}$ is preserved under restriction to dense Zariski-open subsets of the base, as the induced homomorphism between fundamental groups is surjective by [29, Proposition 2.10]. Therefore, after possibly replacing $S$ with $U_{p}$, we may assume that the morphism $p=\nu\circ p^{\nu}\colon S\to T$ is finite and flat. Observe that $p^{\nu}$ is quasi-finite by construction. Moreover, since $\nu$ is separated, the morphism $p^{\nu}$ is proper and consequently finite and surjective.
@@
The restriction to $U_{p}$ is necessary because our argument relies on the “trace operator” for sections, as defined in Equation 24. The key properties of this operator, particularly the one established in Equation 25, are valid when $p$ is flat and finite.
Given a section $\sigma$ of $\mathcal{A}/S$, the proof proceeds by assuming that the group $M_{\sigma}^{\textnormal{rel}}$ is trivial and showing that $\sigma$ must then be unramified. The key idea involves constructing an “unramified family”, i.e. an unramified base change of the universal family, associated with the abelian scheme, which ensures that the cohomology class arising from the logarithm of $\sigma$ “descends”. Applying the Lefschetz hyperplane theorem then imposes strong restrictions on the logarithm of $\sigma$. Combining this with 2.1 and 2.3, we conclude that $\sigma$ is indeed unramified.
In order to construct the aforementioned unramified family, we consider the map $p^{\nu}:S\to T^{\nu}$ and we look at the induced homomorphism on fundamental groups $p^{\nu}_{*}:\pi_{1}(S(\mathbb{C}),s)\to\pi_{1}(T^{\nu}(\mathbb{C}),x)$ where $s\in S(\mathbb{C})$ and $x\in T^{\nu}(\mathbb{C})$ are two base points such that $x=p^{\nu}(s)$. By [31, Corollary 12.19] the inclusion $p^{\nu}_{*}\pi_{1}(S(\mathbb{C}),s)\subseteq\pi_{1}(T^{\nu}(\mathbb{C}),x)$ gives rise to a finite topological cover $q^{\prime}:S^{\prime}(\mathbb{C})\to T^{\nu}(\mathbb{C})$ that satisfies
$q^{\prime}_{*}\pi_{1}(S^{\prime}(\mathbb{C}),s^{\prime})=p^{\nu}_{*}\pi_{1}(S(\mathbb{C}),s),$ (21)
where $s^{\prime}$ is a fixed base point with $q^{\prime}(s^{\prime})=x$. By the Riemann existence theorem (see [26, Exposé XII, Théorème 5.1]) at the level of schemes $q^{\prime}:S^{\prime}\to T^{\nu}$ is a finite étale cover. By the lifting theorem for étale fundamental groups and the relation with the topological fundamental group we can lift $p^{\nu}$ to a morphism $q:S(\mathbb{C})\rightarrow S^{\prime}(\mathbb{C})$ such that $q(s)=s^{\prime}$. Considering the corresponding abelian schemes we obtain the diagram
${\mathcal{A}(\mathbb{C})}$${\mathcal{A}^{\prime}(\mathbb{C})}$${\mathfrak{A}_{g,T^{\nu}}(\mathbb{C})}$${S(\mathbb{C})}$${S^{\prime}(\mathbb{C})}$${T^{\nu}(\mathbb{C}).}$$\scriptstyle{q}$$\scriptstyle{p^{\nu}}$$\scriptstyle{q^{\prime}}$ (22)
Notice that the abelian scheme $\mathcal{A}^{\prime}\to S^{\prime}$ is an unramified family by construction, since $q^{\prime}:S^{\prime}(\mathbb{C})\to T^{\nu}(\mathbb{C})\subseteq\mathbb{A}_{g}(\mathbb{C})$ is in particular an unramified morphism.
@@
We want to point out some algebraic and topological properties of the above construction. By [26, Exposé I, Corollaire 10.2] the finite étale morphism $q^{\prime}$ is obtained as normalization of $T^{\nu}$ in the field extension $\mathbb{C}(S^{\prime})$ of $\mathbb{C}(T^{\nu})$. By universal property of normalization in a field extension, the morphism $q$ is dominant. Since $p$ and $q^{\prime}$ are both finite morphisms then $q$ is quasi-finite and proper, hence a finite surjective morphism as well. Additionally, since $S^{\prime}$ is connected and $T^{\nu}$ is integral and normal it follows that $S^{\prime}$ is normal and irreducible (see [39, Tag 0BQL]).
@@
${\pi_{1}(S(\mathbb{C}),s)}$${\pi_{1}(S^{\prime}(\mathbb{C}),s^{\prime})}$${p^{\nu}_{*}\pi_{1}(S(\mathbb{C}),s),}$$\scriptstyle{q_{*}}$$\scriptstyle{p^{\nu}_{*}}$$\scriptstyle{q^{\prime}_{*}}$ (23)
where $q^{\prime}_{*}:\pi_{1}(S^{\prime}(\mathbb{C}),s^{\prime})\to p^{\nu}_{*}\pi_{1}(S(\mathbb{C}),s)$ is an isomorphism.
@@
In the opposite direction, since $q$ is a finite flat morphism the notion of degree is well-behaved with respect to the ramification index (see [25, Proposition 12.21]). Thus, we can define the trace operator
@@
$\operatorname{Tr}(\tau)(y^{\prime}):=\sum_{y\in q^{-1}(y^{\prime})}m_{y}\tau(y)\in\mathcal{A}^{\prime}_{y^{\prime}},$ (24)
@@
$\operatorname{Tr}(\exp(f))=\exp^{\prime}(\operatorname{Tr}(f))\qquad\textrm{and}\qquad\exp(q^{*}(f^{\prime}))=q^{*}(\exp^{\prime}(f^{\prime})).$ (26)
@@
where $q_{C},q^{\prime}_{C^{\prime}}$ and $p_{C}$ denote the restrictions of the corresponding maps. In this situation, if $C^{\prime}$ and $D$ are still of dimension $1$, we can clearly consider the analogous notions of trace operator and pullback operator as well as the restriction of a section $\sigma:S\to\mathcal{A}$ to $C$, which will be denoted by $\sigma_{C}$. We also have analogous notions of monodromy representations and monodromy groups, e.g. we will write $\rho_{C},\theta_{\sigma_{C}},\textrm{Mon}(\mathcal{A}|_{C})$ and $M^{\textrm{rel}}_{\sigma_{C}}$ for the obvious objects associated to the scheme $\mathcal{A}|_{C}\to C$ and to the section $\sigma_{C}:C\to\mathcal{A}|_{C}$; we do the same for $C^{\prime}$. Moreover, if we denote by $i:C\hookrightarrow S(\mathbb{C})$ and $i^{\prime}:C^{\prime}\hookrightarrow S^{\prime}(\mathbb{C})$ the natural embeddings and choose base points $s\in C$ and $s^{\prime}\in C^{\prime}$, from a topological point of view we obtain the following diagram involving the induced homomorphisms between fundamental groups (for the surjectivity of the upper horizontal arrow see 4.4):
${\pi_{1}(S(\mathbb{C}),s)}$${\pi_{1}(S^{\prime}(\mathbb{C}),s^{\prime})}$${\pi_{1}(C,s)}$${\pi_{1}(C^{\prime},s^{\prime}).}$$\scriptstyle{q_{*}}$$\scriptstyle{{q_{C}}_{\ast}}$$\scriptstyle{i_{*}}$$\scriptstyle{i^{\prime}_{*}}$ (28)
@@ 1. $(i)$
    the morphism $q_{C}:C\to C^{\prime}$ appearing in Equation 27 is finite, the curves $C$ and $C^{\prime}$ are affine and the abelian schemes $\mathcal{A}_{|_{C}}\to C$, $\mathcal{A}^{\prime}_{|_{C^{\prime}}}\to C^{\prime}$ are without fixed part;
@@ 2. $(ii)$
    $\operatorname{Tr}(\sigma_{C})=\operatorname{Tr}(\sigma)|_{C^{\prime}}$ for any $\sigma\in\Sigma^{\textnormal{an}}(S)$ and $\textnormal{Tr}\circ q_{C}^{*}(\tau)=\deg{q}\cdot\tau$ for any $\tau\in\Sigma^{\textnormal{an}}(C^{\prime})$;
@@ 6. $(vi)$
    $\theta_{\sigma_{C}}=\theta_{\sigma}\circ i_{*}$ and $\theta_{\sigma}(\pi_{1}(S(\mathbb{C}),s))=\theta_{\sigma_{C}}(\pi_{1}(C,s))$, for any $\sigma\in\Sigma^{\textnormal{an}}(S)$;
@@ 7. $(vii)$
    if $M_{\sigma}^{\textrm{rel}}=\{0\}$, then $M_{\sigma_{C}}^{\textrm{rel}}=\{0\}$, for any $\sigma\in\Sigma^{\textnormal{an}}(S)$.
@@ 1. $(i)$
    In order to construct the curve $C$ let us consider the general hyperplane section $Y=H\cap S(\mathbb{C})$ and choose a base point $s\in Y$. By Lefschetz theorem on quasi-projective varieties [24, Part II, 1.2, Theorem], the natural map $\pi_{1}(Y,s)\to\pi_{1}(S(\mathbb{C}),s)$ between fundamental groups is surjective. After iterating the process of taking such generic hyperplane sections we obtain an irreducible curve $C_{1}$ with the property that the natural map $\pi_{1}(C_{1},s)\to\pi_{1}(S(\mathbb{C}),s)$ is surjective for a base point $s\in C_{1}$. Note that $C_{1}$ can be assumed to be affine up to remove a finite number of points and using [29, Proposition 2.10]. The restriction of $q$ to $C_{1}$ defines a morphism $q_{C_{1}}$ which is finite: in fact, since every closed embedding is a finite morphism and $q:S\to S^{\prime}$ is a finite morphism (see 4.4), then the restriction $q_{C_{1}}$ is finite. We define $C^{\prime}:=q_{C_{1}}(C_{1})$ and $C:=q^{-1}(C^{\prime})$. Notice that $C^{\prime}$ is an affine curve since $q_{C_{1}}$ is finite and surjective; moreover, $C$ could be no longer irreducible but it is still affine because $q$ is an affine morphism (since it is finite). Furthermore, since $q:S\to S^{\prime}$ is finite then it is closed. We have a closed embedding $i:C\hookrightarrow S(\mathbb{C})$, and we can conclude that the restriction $q_{C}:C\to C^{\prime}$ is a finite morphism. The abelian scheme $\mathcal{A}_{|_{C}}\to C$ is without fixed part by genericity of the hyperplane sections; this implies that also $\mathcal{A}^{\prime}_{|_{C^{\prime}}}\to C^{\prime}$ is without fixed part.
@@ 2. $(ii)$
    This properties are true by construction, since $C=q^{-1}(C^{\prime})$.
@@ 3. $(iii)$
    By $(i)$ we have a surjective homomorphism $\pi_{1}(C_{1},s)\to\pi_{1}(S(\mathbb{C}),s)$. Recall that $C_{1}$ is an irreducible component of $C$. Since the inclusion map $C_{1}\hookrightarrow S(\mathbb{C})$ factors through $C$, then the homomorphism $\pi_{1}(C_{1},s)\to\pi_{1}(S(\mathbb{C}),s)$ factors through $\pi_{1}(C,s)$. Thus, $i_{*}:\pi_{1}(C,s)\to\pi_{1}(S(\mathbb{C}),s)$ is surjective. Then, the surjectivity of $i^{\prime}_{*}:\pi_{1}(C^{\prime},s^{\prime})\to\pi_{1}(S^{\prime}(\mathbb{C}),s^{\prime})$ follows by Equation 28.
@@ 6. $(vi)$
    Since we have the embedding $i:C\hookrightarrow S(\mathbb{C})$, we clearly have $\theta_{\sigma_{C}}(\pi_{1}(C),s)\subseteq\theta_{\sigma}(\pi_{1}(S(\mathbb{C}),s))$. In the other direction, by $(iv)$ it is enough to show that each variation of a logarithm of $\sigma$ induced by an element of $\pi_{1}(S(\mathbb{C}),s)$ can be obtained as a variation of a logarithm of $\sigma_{C}$ induced by the monodromy action of some element of $\pi_{1}(C,s)$; and this is true thanks to the surjectivity of $i_{*}$ which was proven in part $(iii)$. This proves the claim.
@@ 7. $(vii)$
Let us study the exact sequence in cohomology groups induced by Equation 29. By 3.5 we can conclude that no non-zero period can be globally defined on $C$. In other words, we get $H^{0}(C,\Lambda_{C})=0$. In addition, since $C$ is a Stein space we have
@@
By 3.11, we know that the obstruction to define a global logarithm of a section is given by some Galois cohomology class. We want to explain the interplay between the relevant Galois cohomology group and the exact sequence in Equation 30. Let’s fix the notation $G:=\pi_{1}(C,s)$ for a base point $s\in C$ and denote by $\rho_{C}:G\to\textrm{Mon}(\mathcal{A}|_{C})\subseteq\textrm{Sp}_{2g}(\mathbb{Z})$ the monodromy representation associated to the abelian scheme $\mathcal{A}|_{C}\to C$. We give a description of the group $H^{1}(C,\Lambda_{C})$ in terms of the Galois cohomology with respect to the action of $G$ on $\mathbb{Z}^{2g}$ induced by the projection $\rho_{C}$. We have the following
@@
It follows from [1, Proposition 2.1.1] since the universal cover of $C$ is contractible. One can also generalize the independent proof given in [13, Proposition 4.1]. ∎
Therefore, thanks to the exact sequence Equation 30 and 4.6 we obtain a surjective group homomorphism
@@
Notice that by 4.5 also the curve $C^{\prime}$ is a Stein space. Moreover, by 3.5 we still obtain $H^{0}(C^{\prime},\Lambda_{C^{\prime}})=0$. Therefore, the analogous results of Equation 30 and 4.6 hold if we replace the curve $C$ by the curve $C^{\prime}$. In particular, for some base point $s^{\prime}\in C^{\prime}$ we have the surjective group homomorphism